OpenStudy (mizuki):

You toss a coin and roll a number cube. Find P(heads and an even number). 1 1/4 1/6 1/12

OpenStudy (mizuki):

B?

OpenStudy (kittiwitti1):

A coin is double sided so[\text{Coin chances}\rightarrow\frac{1}{2}\]A number cube has 6 faces/sides so\[\text{Number cube chances}\rightarrow\frac{3}{6}\]

OpenStudy (kittiwitti1):

Oops. \[\text{Coin chances}\rightarrow\frac{1}{2}\]

OpenStudy (mizuki):

C?

OpenStudy (kittiwitti1):

I forgot where to go from here T_T